# Representing Anatomical Trees by Denoising Diffusion of Implicit Neural Fields

### v1

@@ 1 Introduction
Recently, INRs (or neural fields, implicit fields) have been proposed to address the shortcomings of explicit representations, which use multi-layer perceptrons (MLP) to fit a continuous function that implicitly represents a signal of interest [21, 23, 29] as level sets.
This approach not only learns a representation of arbitrary shapes with a small number of parameters while still achieving a high surface accuracy, but also supports reconstructions at any arbitrary resolution. To this end, we propose to use INRs for representing anatomical trees and demonstrate their efficiency and effectiveness in modeling simple and complex, vasculature and airways, in 2D and 3D, on both synthetic and real data.
@@ 2 Methodology
Similar to [21, 22], our MLP architecture is a fully-connected network with $L$ layers, of hidden size $D$, and ReLU activation functions. To learn high-fidelity neural representations of anatomical trees, we follow [21] to first instantiate all training instances $S_{i}$ in Lipschitz domain $\Omega\in[-1,1]$ and then adaptively sample points and ground-truth occupancies both inside and outside the 3D surface of $S_{i}$, to supervise the optimization of each MLP until convergence. LABEL:fig:optimization illustrates this optimization process.
@@ 3 Datasets, Results, and Implementation Details
Figure 3: Compression vs Reconstruction Accuracy: (Top) We decimate/downscale the mesh/volume to occupy $\approx$ same memory space as INR, and report the reconstruction error using Chamfer distance (CD) and edge length loss (Edge) for meshes and INRs, and $L_{1}$ and $L_{2}$ for volumes. Note the higher error for meshes and volumes w.r.t INRs for the same storage. (Bottom) Illustration of GT and downsampled volumes. Notice the disconnected components.
@@
[…]
Figure 6: Tree Statistics computed on VascuSynth. (a) Distance between INRs of trees with similar bifurcations is low and increases as the complexity increases; (b) t-SNE plot of the space of trees as INRs; (c-d) Histograms of tortuosity and total length for ground truth and generated trees.
[…]
@@
Once trees are represented via INRs, we study the space of INRs. The t-SNE plot in Figure 5(a) illustrates how trees with similar number of bifurcations, i.e. topological complexities, tend to cluster in the 2D t-SNE space. However, we observe cases where trees with the same number of branches are distant, which may be due to how t-SNE projects high-dimensional samples to 2D thus distort inter-sample distances. Therefore, to further investigate, we compute the matrix of $L_{2}$ distances between INRs of VascuSynth trees with $i$ and $j$ bifurcations, where $i,j\in\{1,\dots,12\}$ (Fig. 5(b)). The reported distances are the average across $10$ trees per bifurcation count. We notice low ($\approx 0$) values along the diagonal, indicating that trees with the same branch count have similar INRs. As the difference in branch count increases, the distance increases, visible by the color gradient in Fig. 5(b). Furthermore, for a constant difference in branch count (e.g., $|3-4|=1=|8-9|$), the dissimilarity is more pronounced with larger branch counts, despite the numerical difference being the same.
@@
Training a DDM on INR-based trees results in a model that captures the tree distribution and can be sampled using DDIM [31] to generate novel trees. First, we show qualitative results of novel INR-based trees generated by our model in Figure 8. Next, since the evaluation of unconditional generation of tree-like structures can be challenging due to lack of direct correspondence to ground truth data, we follow [4, 5] to quantitatively assess the samples generated using diffusion. Specifically, we use minimum matching distance (MMD), coverage (COV), and 1-nearest neighbor accuracy (1-NNA) to measure quality, diversity, and plausibility, respectively. Moreover, since the generations are random, we run the evaluation 3 times and report the mean values in Table 3. Additionally, we follow [6, 37] and report vessel-based metrics in Figs. 5(c) and 5(d), where we see high similarities in tortuosity and centerline length histograms between ground truth training data and synthesized trees.

### v2

@@ 1 Introduction
Recently, INRs (or neural fields, implicit fields) have been proposed to address the shortcomings of explicit representations, which use multi-layer perceptrons (MLP) to fit a continuous function that implicitly represents a signal of interest [21, 23, 29] as level sets. This approach not only learns a representation of arbitrary shapes with a small number of parameters while still achieving a high surface accuracy, but also supports reconstructions at any arbitrary resolution. To this end, we propose to use INRs for representing anatomical trees and demonstrate their efficiency and effectiveness in modeling simple and complex, vasculature and airways, in 2D and 3D, on both synthetic and real data.
@@ 2 Methodology
Similar to [21, 22], our MLP architecture is a fully-connected network with $L$ layers, of hidden size $D$, and ReLU activation functions. To learn high-fidelity neural representations of anatomical trees, we follow [21] to first instantiate all training instances $S_{i}$ in Lipschitz domain $\Omega\in[-1,1]$ and then adaptively sample points and ground-truth occupancies both inside and outside the 3D surface of $S_{i}$, to supervise the optimization of each MLP until convergence. Figure 3 illustrates this optimization process.
@@ 3 Datasets, Results, and Implementation Details
Figure 4: Compression vs Reconstruction Accuracy: (a) We decimate/downscale the mesh/volume to occupy $\approx$ same memory space as INR, and report the reconstruction error using Chamfer distance (CD) and edge length loss (Edge) for meshes and INRs, and $L_{1}$ and $L_{2}$ for volumes. Note the higher error for meshes and volumes w.r.t INRs for the same storage. (b) Illustration of GT and downsampled volumes. Notice the disconnected components.
@@
Figure 6: Arbitrary Resolution: (a) Comparison of 2x, 4x, and 8x zoom on an IntRA sample represented as a volumetric grid (top) and INR (bottom). The resolution for each sub-figure is shown on top as volume3/INR3. Notice the smoothness of the surface even at 8x zoom. (b) Zoomed-in regions of a VascuSynth mesh reconstructed from INRs and ground truth at different mesh resolutions displayed using faces and edges.
Figure 7: Tree Statistics computed on VascuSynth. (a) Distance between INRs of trees with similar bifurcations is low and increases as the complexity increases; (b) t-SNE plot of the space of trees as INRs; (c-d) Histograms of tortuosity, total length and average radius for ground truth and generated trees.
@@
Once trees are represented via INRs, we study the space of INRs. The t-SNE plot in Figure 7(b) illustrates how trees with similar number of bifurcations, i.e. topological complexities, tend to cluster in the 2D t-SNE space. However, we observe cases where trees with the same number of branches are distant, which may be due to how t-SNE projects high-dimensional samples to 2D thus distort inter-sample distances. Therefore, to further investigate, we compute the matrix of $L_{2}$ distances between INRs of VascuSynth trees with $i$ and $j$ bifurcations, where $i,j\in\{1,\dots,12\}$ (Fig. 7(a)). The reported distances are the average across $10$ trees per bifurcation count. We notice low ($\approx 0$) values along the diagonal, indicating that trees with the same branch count have similar INRs. As the difference in branch count increases, the distance increases, visible by the color gradient in Fig. 7(a). Furthermore, for a constant difference in branch count (e.g., $|3-4|=1=|8-9|$), the dissimilarity is more pronounced with larger branch counts, despite the numerical difference being the same.
@@
Training a DDM on INR-based trees results in a model that captures the tree distribution and can be sampled using DDIM [31] to generate novel trees. First, we show qualitative results of novel INR-based trees generated by our model in Figure 9. Next, since the evaluation of unconditional generation of tree-like structures can be challenging due to lack of direct correspondence to ground truth data, we follow [4, 5] to quantitatively assess the samples generated using diffusion. Specifically, we use minimum matching distance (MMD), coverage (COV), and 1-nearest neighbor accuracy (1-NNA) to measure quality, diversity, and plausibility, respectively. Moreover, since the generations are random, we run the evaluation 3 times and report the mean values in Table 3. Additionally, we follow [6, 37] and report vessel-based metrics in Figs. 7(c) and 7(d), where we see high similarities in tortuosity and centerline length histograms between ground truth training data and synthesized trees.
@@ 4 Conclusion
5 Acknowledgements
The authors are grateful to Paula Feldman for the provision of evaluation scripts used for computing tortuosity and other tree-specific metrics. This research was enabled in part by support provided by the Natural Sciences and Engineering Research Council of Canada (NSERC), and the computational resources provided by Digital Research Alliance of Canada and NVIDIA Corporation.